# Continuous difference-in-differences with double/debiased machine learning

### v5

@@ Abstract
This paper extends difference-in-differences to settings with continuous treatments. Specifically, the average treatment effect on the treated (ATT) at any level of treatment intensity is identified under a conditional parallel trends assumption. Estimating the ATT in this framework requires first estimating infinite-dimensional nuisance parameters, particularly the conditional density of the continuous treatment, which can introduce substantial bias. To address this challenge, we propose estimators for the causal parameters under the double/debiased machine learning framework and establish their asymptotic normality. Additionally, we provide consistent variance estimators and construct uniform confidence bands based on a multiplier bootstrap procedure. To demonstrate the effectiveness of our approach, we apply our estimators to the 1983 Medicare Prospective Payment System (PPS) reform studied by Acemoglu and Finkelstein (2008), reframing it as a DiD with continuous treatment and nonparametrically estimating its effects.
@@ 1 Introduction
Despite its widespread use in empirical research, the theoretical foundation for continuous DiD remains relatively underdeveloped, particularly in comparison to the extensive body of literature on DiD with binary or discrete treatments (see Roth et al. (2023), de Chaisemartin and D’Haultfoeuille (2023), Callaway (2023) for recent overviews). A few recent studies have begun to bridge this gap, notably de Chaisemartin et al. (2022), D’Haultfoeuille et al. (2023), and Callaway et al. (2024). For instance, D’Haultfoeuille et al. (2023) extend the change-in-changes model of Athey and Imbens (2006) to accommodate continuous treatments, while de Chaisemartin et al. (2022) examine the average slope of stayers in the continuous DiD setting. Our paper is closely related to Callaway et al. (2024), which studies continuous DiD in the commonly used two-way fixed effect (TWFE) regression framework. Callaway et al. (2024) demonstrate that, under TWFE, the regression parameter of interest can be decomposed as weighted integrals of either the average treatment parameters across treatment intensities with potentially negative weights or average causal responses with selection bias but nonnegative weights. They also provide data-driven nonparametric estimators for these causal parameters that are rate optimal.
In this paper, we focus on the average treatment effect on the treated (ATT) for any given continuous treatment intensity. Although this parameter is one of several investigated in Callaway et al. (2024), our primary contribution is to incorporate covariates nonparametrically into both the identification and estimation procedures. Specifically, we modify the parallel trends assumption in Callaway et al. (2024) by conditioning on covariates in a manner analogous to the “conditional parallel trends” assumption used in DiD for binary or discrete treatments; see Heckman et al. (1997, 1998), Abadie (2005), Chang (2020), and Sant’Anna and Zhao (2020), for example. As noted in Abadie (2005), an unconditional parallel trends assumption can be restrictive if covariates that influence outcome dynamics have different distributions across treatment and control groups. By conditioning on such covariates, we obtain a more robust framework for identifying and estimating the ATT in continuous treatment settings.
@@
To illustrate the usefulness of our method, we revisit Acemoglu and Finkelstein (2008), which examines the impact of the 1983 Medicare payment system (PPS) reform on the healthcare industry. Since the PPS reform affected hospitals with varying proportions of Medicare inpatients differently, the share of Medicare inpatients can be interpreted as a continuous treatment variable. This makes Acemoglu and Finkelstein (2008) an exemplary case for applying our methods. Thus, we nonparametrically estimate the ATTs of the PPS reform in a continuous DiD context, providing a more detailed understanding of the effects of this policy reform. In particular, contrasting with the linear estimates from Acemoglu and Finkelstein (2008), our results suggest significant heterogeneity in the impact of the PPS reform across hospitals with different shares of Medicare inpatients.
@@ 3 Orthogonal Scores
where $\Delta Y:=Y_{t}-Y_{t-1}$. First, given the continuous nature of the treatment intensity, $\theta_{0}(d)$ cannot be estimated nonparametrically at root-$N$ rate. This relates to a class of non-regular parameters involving continuous treatment variables; see Galvao and Wang (2015), Kennedy et al. (2017), Su et al. (2019), Semenova and Chernozhukov (2021), Fan et al. (2022), and Colangelo and Lee (2025) for example. Moreover, a score based on the above expression does not satisfy Neyman orthogonality, and an adjustment term has to be added.
@@ 4 Estimation and Inference
Next, we outline the regularity conditions required to establish the asymptotic properties of our DML estimators. We focus on the panel case and present the analogous results for the repeated cross-sections in Appendix B. For notation simplicity, let $\mathcal{D}:=(d_{L},d_{H})$, and let $\mathcal{X}$ and $\Delta\mathcal{Y}$ denote the support of $X$ and $\Delta Y$ respectively.
@@
The kernel function is central to our analysis. In addition to its well-established theoretical properties for estimating the density $f_{D}(d)$, we also use it to approximate the point mass at $D=d$ and the conditional density $f_{D|X}(d|X)$. Assumption 4.6 imposes the standard regularity conditions on the kernel function, which are essential for establishing the asymptotic normality of our estimator. Assumption 4.7 requires smoothness and boundedness of the outcome variable and relevant distributions, while Assumption 4.8 specifies conditions on the kernel bandwidth and the quality of the nonparametric nuisance estimators.
@@
Assumption 4.8 (a) and (b) give $h=o(N^{-1/5})$ and $h=\omega(\varepsilon_{N}^{2})$. However, consistency of our variance estimator additionally requires $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$, which imposes a more restrictive lower bound on $h$. Moreover, whereas the standard DML literature assumes the nuisance estimators to converge at rate $\varepsilon_{N}=o(N^{-1/4})$, we allow the conditional density $\hat{f}_{h}$ to converge at a slower rate $h^{-1/2}\varepsilon_{N}$. This relaxation does not contradict the existing DML results for regular parameters; our target parameter is non-regular and cannot be estimated nonparametrically at $\sqrt{N}$ rate because of the continuous treatment. Finally, although our results assume a deterministic kernel bandwidth, they should extend to data-driven choices (for example, the adaptive procedure in Bibaut and van der Laan (2017)), which we leave to future work.
@@
Suppose assumptions 2.1, 2.3, 2.4, 2.5, 4.6, 4.7, and 4.8 hold. Then, for $d\in\mathcal{D}$, if $\epsilon_{N}=o(N^{-1/4})$,
@@
The proof builds on the DML framework of Chernozhukov et al. (2018), modified to accommodate kernel smoothing. The asymptotic variance has two components, both depending inversely on the kernel bandwidth $h$: one arising from the kernels in the orthogonal score $\psi_{h}$, and the other from the linear expansion of the estimator with respect to the kernel density estimator $\hat{f}_{D}(d)$. Since $h$ is a function of the sample size $N$ under our assumptions, we index the asymptotic variance by $N$ to reflect this dependence. Therefore, our estimator $\widehat{ATT}(d)$ attains a convergence rate of $\sqrt{Nh}$, which, though slower than the parametric rate $\sqrt{N}$, is comparable to the optimal rate for one-dimensional nonparametric regression estimation.
@@
Suppose the conditions of Theorem 4.2 hold and assume that $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$. Then
@@
This assumption differs from the pointwise case in two key ways. First, we require that, uniformly over the compact domain $\mathcal{D}$, the nuisance estimator $\hat{f}_{h}(d|X)$ remains bounded and has rate $h^{-1/2}\epsilon_{N}$. Second, we assume that the estimated density and conditional density to have bounded derivatives with probability tending to $1$, ensuring that the score functions are Lipschitz continuous on $\mathcal{D}$. These additional assumptions are mild and can be enforced during estimation procedures. With these modified assumptions, the linear expansion of the bootstrap estimators holds uniformly over $d\in\mathcal{D}$.
@@
Suppose assumptions 2.1, 2.3, 2.4, 2.5, 4.6, 4.7, 4.9, and 4.10 hold. Then for $\epsilon_{N}=o(N^{-1/4})$
@@ 5 Simulation
In our simulations, the nuisance parameters $P(D=0|X)$, $f_{h}(d|X)=E[K_{h}(D-d)|X]$, $E[Y_{t}-Y_{t-1}|X,D=0]$, and $E\big[\frac{T-\lambda}{\lambda(1-\lambda)}Y\big|X,D=0\big]$ are estimated nonparametrically using random forests each with 200 trees of maximum depth 20. Throughout our simulations, we also use an undersmoothing kernel bandwidth $h=1.06\hat{\sigma}_{\tilde{D}}N^{-1/4}$, where $\hat{\sigma}_{\tilde{D}}$ is the estimated standard deviation of positive treatment intensities. We consider sample sizes $N=2000$ and $10000$ for both panel and repeated cross-sectional settings, and we conduct $B=500$ simulations in each setting. The DGP implies the true $ATT(d)=-0.5d^{2}$, and we focus on a specific treatment intensity $d=0.9$. Notably, the continuous treatment variable is dependent on the correlated high-dimensional covariates in a nonlinear way. Additionally, the DGPs suggest that the effective sample size should be small at the target intensity, which adds another layer of difficulty for estimation.
@@ 7 Conclusion
This paper studies difference-in-differences models with continuous treatments. Our identification results are based on a conditional parallel trends assumption, allowing researchers to account for covariates nonparametrically. Under the double/debiased machine learning framework, we develop nonparametric estimators for the average treatment effect on the treated at each continuous treatment intensity and establish their asymptotic properties. Monte Carlo simulations demonstrate that our estimators perform well despite the highly non-linear relationship between the continuous treatment and the high-dimensional covariates. To demonstrate the empirical relevance of our methodology, we re-examine the research questions posed in Acemoglu and Finkelstein (2008) by applying our estimators to their dataset and obtaining new empirical insights. The extension of difference-in-differences models to the continuous treatment setting has important implications for empirical research. Our methods provide researchers with new tools for examining the impacts of continuous treatment variables.
@@
In this section, we present the asymptotic theory for the repeated cross-sectional setting, following a structure analogous to the panel case discussed in detail in the main text. For notation simplicity, let $\mathcal{D}=(d_{L},d_{H})$ denote the interior of the support of $D>0$. Let $\mathcal{X}$ and $\mathcal{Y}^{\lambda}$ denote the support of $X$ and $Y^{\lambda}$ respectively. We begin by introducing regularity conditions that allow us to establish the asymptotic normality of our estimator and the consistency of the corresponding variance estimator.
@@
Suppose Assumptions 2.2, 2.3, 2.4, 2.5, 4.6, 7.11, and 7.12 hold. Then, for $d\in\mathcal{D}$, if $\epsilon_{N}=o(N^{-1/4})$,
@@
If the conditions in Theorem 7.5 hold and assume $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$, then
@@
Suppose Assumptions 2.2, 2.3, 2.4, 2.5, 4.6, 7.11, and 7.13 hold. Then for $\epsilon_{N}=o(N^{-1/4})$
@@
Then, let $\delta>0$ be given, as $n\to\infty$,
$P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\delta(Nh)^{-1/2}\Big)$ (C.197)
$\leq$ $\epsilon+P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\mathbf{1}\{A_{N}(\epsilon)\}\geq\delta(Nh)^{-1/2}\Big)$ (C.198)
$\leq$ $\epsilon+E\Big[P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\delta(Nh)^{-1/2}|I_{k}^{c}\Big)\mathbf{1}\{A_{N}(\epsilon)\}\Big]$ (C.199)
$\leq$ $\epsilon+E\Bigg[\frac{E[\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}|I_{k}^{c}]}{\delta(Nh)^{-1/2}}\mathbf{1}\{A_{N}(\epsilon)\}\Bigg]$ (C.200)
$\lesssim$ $\epsilon+\frac{\sqrt{\frac{h^{-1}\epsilon_{N}^{2}\log(N)}{N}}+\frac{\log^{\frac{3}{2}}(N)h^{-1}}{N}}{\delta(Nh)^{-1/2}}$ (C.201)

### v6

@@ Abstract
This paper extends difference-in-differences to settings with continuous treatments. Specifically, the average treatment effect on the treated (ATT) at any level of treatment intensity is identified under a conditional parallel trends assumption. Estimating the ATT in this framework requires first estimating infinite-dimensional nuisance parameters, particularly the conditional density of the continuous treatment, which can introduce substantial bias. To address this challenge, we propose estimators for the causal parameters under the double/debiased machine learning framework and establish their asymptotic normality. Additionally, we provide consistent variance estimators and construct uniform confidence bands based on a multiplier bootstrap procedure. To demonstrate the effectiveness of our approach, we revisit a previous study on the 1983 Medicare Prospective Payment System reform, reframing it as a DiD with continuous treatment and non-parametrically estimating its effects.
@@ 1 Introduction
Despite its widespread use in empirical research, the theoretical foundation for continuous DiD remains relatively underdeveloped, particularly in comparison to the extensive body of literature on DiD with binary or discrete treatments (see Roth et al. (2023), de Chaisemartin and D’Haultfoeuille (2023), Callaway (2023) for recent overviews). A few recent studies have begun to bridge this gap, notably de Chaisemartin et al. (2022), D’Haultfoeuille et al. (2023), and Callaway et al. (2024). For instance, D’Haultfoeuille et al. (2023) extend the change-in-changes model of Athey and Imbens (2006) to accommodate continuous treatments, while de Chaisemartin et al. (2022) examine the average slope of stayers in the continuous DiD setting. Our paper is closely related to Callaway et al. (2024), which studies continuous DiD in the commonly used two-way fixed effect (TWFE) regression framework. Callaway et al. (2024) demonstrate that, under TWFE, the regression parameter of interest can be decomposed as weighted integrals of either the average treatment parameters across treatment intensities with potentially negative weights or average causal responses with selection bias but nonnegative weights. They also provide data-driven non-parametric estimators for these causal parameters that are rate optimal.
In this paper, we focus on the average treatment effect on the treated (ATT) for any given continuous treatment intensity. Although this parameter is one of several investigated in Callaway et al. (2024), our primary contribution is to incorporate covariates non-parametrically into both the identification and estimation procedures. Specifically, we modify the parallel trends assumption in Callaway et al. (2024) by conditioning on covariates in a manner analogous to the “conditional parallel trends” assumption used in DiD for binary or discrete treatments; see Heckman et al. (1997, 1998), Abadie (2005), Chang (2020), and Sant’Anna and Zhao (2020), for example. As noted in Abadie (2005), an unconditional parallel trends assumption can be restrictive if covariates that influence outcome dynamics have different distributions across treatment and control groups. By conditioning on such covariates, we obtain a more robust framework for identifying and estimating the ATT in continuous treatment settings.
@@
To illustrate the usefulness of our method, we revisit Acemoglu and Finkelstein (2008), which examines the impact of the 1983 Medicare payment system (PPS) reform on the healthcare industry. Since the PPS reform affected hospitals with varying proportions of Medicare inpatients differently, the share of Medicare inpatients can be interpreted as a continuous treatment variable. This makes Acemoglu and Finkelstein (2008) an exemplary case for applying our methods. Thus, we non-parametrically estimate the ATTs of the PPS reform in a continuous DiD context, providing a more detailed understanding of the effects of this policy reform. In particular, contrasting with the linear estimates from Acemoglu and Finkelstein (2008), our results suggest significant heterogeneity in the impact of the PPS reform across hospitals with different shares of Medicare inpatients.
@@ 3 Orthogonal Scores
where $\Delta Y:=Y_{t}-Y_{t-1}$. First, given the continuous nature of the treatment intensity, $\theta_{0}(d)$ cannot be estimated non-parametrically at root-$N$ rate. This relates to a class of non-regular parameters involving continuous treatment variables; see Galvao and Wang (2015), Kennedy et al. (2017), Su et al. (2019), Semenova and Chernozhukov (2021), Fan et al. (2022), and Colangelo and Lee (2025) for example. Moreover, a score based on the above expression does not satisfy Neyman orthogonality, and an adjustment term has to be added.
@@ 4 Estimation and Inference
Next, we outline the regularity conditions required to establish the asymptotic properties of our DML estimators. We focus on the panel case and present the analogous results for the repeated cross-sections in Appendix B. For notational simplicity, let $\mathcal{D}$ denote a closed sub-interval of $(d_{L},d_{H})$ whose boundary points can be chosen arbitrarily close to $d_{L}$ and $d_{H}$, and let $\mathcal{X}$ and $\Delta\mathcal{Y}$ denote the supports of $X$ and $\Delta Y$, respectively.
@@
The kernel function is central to our analysis. In addition to its well-established theoretical properties for estimating the density $f_{D}(d)$, we also use it to approximate the point mass at $D=d$ and the conditional density $f_{D|X}(d|X)$. Assumption 4.6 imposes the standard regularity conditions on the kernel function, which are essential for establishing the asymptotic normality of our estimator. Assumption 4.7 requires smoothness and boundedness of the outcome variable and relevant distributions, while Assumption 4.8 specifies conditions on the kernel bandwidth and the quality of the non-parametric nuisance estimators.
@@
Assumption 4.8 (a) and (b) give $h=o(N^{-1/5})$ and $h=\omega(\varepsilon_{N}^{2})$. However, consistency of our variance estimator additionally requires $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$, which imposes a more restrictive lower bound on $h$. Moreover, whereas the standard DML literature assumes the nuisance estimators to converge at rate $\varepsilon_{N}=o(N^{-1/4})$, we allow the conditional density $\hat{f}_{h}$ to converge at a slower rate $h^{-1/2}\varepsilon_{N}$. This relaxation does not contradict the existing DML results for regular parameters; our target parameter is non-regular and cannot be estimated non-parametrically at $\sqrt{N}$ rate because of the continuous treatment. Finally, although our results assume a deterministic kernel bandwidth, they should extend to data-driven choices (for example, the adaptive procedure in Bibaut and van der Laan (2017)), which we leave to future work.
@@
Suppose assumptions 2.1, 2.3, 2.4, 2.5, 4.6, 4.7, and 4.8 hold. Then, for $d\in\mathcal{D}$, if $\varepsilon_{N}=o(N^{-1/4})$,
@@
The proof builds on the DML framework of Chernozhukov et al. (2018), modified to accommodate kernel smoothing. The asymptotic variance has two components, both depending inversely on the kernel bandwidth $h$: one arising from the kernels in the orthogonal score $\psi_{h}$, and the other from the linear expansion of the estimator with respect to the kernel density estimator $\hat{f}_{D}(d)$. Since $h$ is a function of the sample size $N$ under our assumptions, we index the asymptotic variance by $N$ to reflect this dependence. Therefore, our estimator $\widehat{ATT}(d)$ attains a convergence rate of $\sqrt{Nh}$, which, though slower than the parametric rate $\sqrt{N}$, is comparable to the optimal rate for one-dimensional non-parametric regression estimation.
@@
Suppose the conditions of Theorem 4.2 hold and assume that $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$. Then, for $d\in\mathcal{D}$,
@@
This assumption differs from the pointwise case in two key ways. First, we require that, uniformly over $\mathcal{D}$, the nuisance estimator $\hat{f}_{h}(d|X)$ remains bounded and has rate $h^{-1/2}\varepsilon_{N}$. Second, we assume that the estimated density and conditional density to have bounded derivatives with probability tending to $1$, ensuring that the score functions are Lipschitz continuous on $\mathcal{D}$. These additional assumptions are mild and can be enforced during estimation procedures. With these modified assumptions, the linear expansion of the bootstrap estimators holds uniformly over $d\in\mathcal{D}$.
@@
Suppose assumptions 2.1, 2.3, 2.4, 2.5, 4.6, 4.7, 4.9, and 4.10 hold. Then, for $d\in\mathcal{D}$, if $\varepsilon_{N}=o(N^{-1/4})$,
@@ 5 Simulation
In our simulations, the nuisance parameters $P(D=0|X)$, $f_{h}(d|X)=E[K_{h}(D-d)|X]$, $E[Y_{t}-Y_{t-1}|X,D=0]$, and $E\big[\frac{T-\lambda}{\lambda(1-\lambda)}Y\big|X,D=0\big]$ are estimated non-parametrically using random forests each with 200 trees of maximum depth 20. Throughout our simulations, we also use an undersmoothing kernel bandwidth $h=1.06\hat{\sigma}_{\tilde{D}}N^{-1/4}$, where $\hat{\sigma}_{\tilde{D}}$ is the estimated standard deviation of positive treatment intensities. We consider sample sizes $N=2000$ and $10000$ for both panel and repeated cross-sectional settings, and we conduct $B=500$ simulations in each setting. The DGP implies the true $ATT(d)=-0.5d^{2}$, and we focus on a specific treatment intensity $d=0.9$. Notably, the continuous treatment variable is dependent on the correlated high-dimensional covariates in a nonlinear way. Additionally, the DGPs suggest that the effective sample size should be small at the target intensity, which adds another layer of difficulty for estimation.
@@ 7 Conclusion
This paper studies difference-in-differences models with continuous treatments. Our identification results are based on a conditional parallel trends assumption, allowing researchers to account for covariates non-parametrically. Under the double/debiased machine learning framework, we develop non-parametric estimators for the average treatment effect on the treated at each continuous treatment intensity and establish their asymptotic properties. Monte Carlo simulations demonstrate that our estimators perform well despite the highly non-linear relationship between the continuous treatment and the high-dimensional covariates. To demonstrate the empirical relevance of our methodology, we re-examine the research questions posed in Acemoglu and Finkelstein (2008) by applying our estimators to their dataset and obtaining new empirical insights. The extension of difference-in-differences models to the continuous treatment setting has important implications for empirical research. Our methods provide researchers with new tools for examining the impacts of continuous treatment variables.
@@
In this section, we present the asymptotic theory for the repeated cross-sectional setting, following a structure analogous to the panel case discussed in detail in the main text. For notational simplicity, let $\mathcal{D}$ denote a closed sub-interval of $(d_{L},d_{H})$ whose boundary points can be chosen arbitrarily close to $d_{L}$ and $d_{H}$, and let $\mathcal{X}$ and $\mathcal{Y}^{\lambda}$ denote the supports of $X$ and $Y^{\lambda}$, respectively. We begin by introducing regularity conditions that allow us to establish the asymptotic normality of our estimator and the consistency of the corresponding variance estimator.
@@
Suppose Assumptions 2.2, 2.3, 2.4, 2.5, 4.6, 7.11, and 7.12 hold. Then, for $d\in\mathcal{D}$, if $\varepsilon_{N}=o(N^{-1/4})$,
@@
If the conditions in Theorem 7.5 hold and assume $h^{-2}\varepsilon_{N}^{2}+h^{-3}N^{-1}=o(1)$, then, for $d\in\mathcal{D}$,
@@
Suppose Assumptions 2.2, 2.3, 2.4, 2.5, 4.6, 7.11, and 7.13 hold. Then, for $d\in\mathcal{D}$, if $\varepsilon_{N}=o(N^{-1/4})$,
@@
Then, let $\zeta>0$ be given, as $N\to\infty$,
$P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\zeta(Nh)^{-1/2}\Big)$ (C.197)
$=$ $E\Big[P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\zeta(Nh)^{-1/2}|I_{k}^{c}\Big)\Big]$ (C.198)
$=$ $E\Big[P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\zeta(Nh)^{-1/2}|I_{k}^{c}\Big)\mathbf{1}\{A_{N}^{c}(\epsilon)\}\Big]$ (C.199)
$+E\Big[P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\zeta(Nh)^{-1/2}|I_{k}^{c}\Big)\mathbf{1}\{A_{N}(\epsilon)\}\Big]$ (C.200)
$\leq$ $\epsilon+E\Big[P\Big(\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}\geq\zeta(Nh)^{-1/2}|I_{k}^{c}\Big)\mathbf{1}\{A_{N}(\epsilon)\}\Big]$ (C.201)
$\leq$ $\epsilon+E\Bigg[\frac{E[\|\dot{P}_{n,k}^{*}[\Delta(\hat{\eta}_{k}(d),d)]\|_{\mathcal{F_{N}}}|I_{k}^{c}]}{\zeta(Nh)^{-1/2}}\mathbf{1}\{A_{N}(\epsilon)\}\Bigg]$ (C.202)
$\lesssim$ $\epsilon+\frac{\sqrt{\frac{h^{-1}\epsilon_{N}^{2}\log(N)}{N}}+\frac{\log^{\frac{3}{2}}(N)h^{-1}}{N}}{\zeta(Nh)^{-1/2}}$ (C.203)